The 2021 balance sheet for Hallwood Industries, Inc., is shown below.

HALLWOOD INDUSTRIES, INC.
Balance Sheet
December 31, 2021
($ in thousands)
Assets
Cash $ 360
Short-term investments 310
Accounts receivable 360
Inventory 370
Property, plant, and equipment (net) 2,600
Total assets $ 4,000
Liabilities and Shareholders’ Equity
Current liabilities $ 560
Long-term liabilities 510
Paid-in capital 1,550
Retained earnings 1,380
Total liabilities and shareholders’ equity $ 4,000


The company’s 2021 income statement reported the following amounts ($ in thousands):

Net sales $ 6,200
Interest expense 50
Income tax expense 150
Net income 320


Required:
1. Calculate the current ratio. (Round your answer to 2 decimal places.)
2. Calculate the acid-test ratio. (Round your answer to 3 decimal places.)
3. Calculate the debt to equity ratio. (Round your answer to 2 decimal places.)
4. Calculate the times interest earned ratio. (Round your answer to 1 decimal place.)

1. Current ratio
2. Acid-test ratio
3. Debt to equity ratio
4. Times interest earned ratio times
Answer #1

Current Ratio = Current assets/Current liabilities  

= (360+310+360+370)/560

= 2.50

Acid test ratio = Quick assets/Current liabilities

= (4000-2600-370)/560

= 1.84

Debt to equity ratio = Total Liabilities/Total stockholders equity

= (560+510)/(1550+1380)

= 0.37

Times interest earned ratio = Net income+Interest expense+Tax expenses/Interest expense

= (320+150+50)/50

= 10.4